OpenStudy (shainabeans):

find sinx, given cos2x=3/5 and x terminates in quad II

OpenStudy (dumbcow):

use trig identity : cos2x = 1-2sin^2 --> 3/5 = 1-2sin^2 --> sin^2 = 1/5 --> sin x = sqrt5/5